The length of a rectangle is 8 feet longer than its wide. If each side is increased 8 feet, then the area is multiplied by 3. What was the size of the original rectangle?
The width or short side is
The Length or the long side is

Answer by jorel555(1290)

Let w=width of first rectangle, and a=area. Original length, l, is w+8;
w(w+8)=a
w^2+8w=a
3w^2+24w=3a
The second rectangle is;
(w+8)(w+16)=3a
w^2+24w+128=3w^2+24w
2w^2=128
w^2=64
w=8
The original rectangle was 8 ft in width; 16 feet length!!!!!!!!!!